# Detecting Flaky Tests in Quantum Software: A Dynamic Approach

### v1

@@ 4 Study Results
Table 2: Test suite size and flakiness across terra releases. For each release, we report the total number of test cases, the number of non-skipped tests, the number of flaky tests (i.e., those that both pass and fail across 10,000 runs), the flakiness rate, and the contribution of that release to the global set of unique flaky tests. The flakiness rate is computed as the ratio of flaky test count over total test count. The share of all unique flaky tests is computed as the ratio of the flaky test count for a given release over the total number of unique flaky tests across all releases ($n=290$). The final two rows aggregate totals across all releases.
@@ 4.3.2 Theoretical analysis
Figure 2: Wilson confidence intervals for the true failure probability $p$ as a function of the empirical estimate $\hat{p}$. For each confidence level (90%, 95%, 99%), the upper and lower bounds are shown as paired lines with matching color and line type. The reference line $p=\hat{p}$ is shown with a dot-dash pattern.
Figure 2 shows how the Wilson interval varies with $\hat{p}$ at different confidence levels. As $\hat{p}$ decreases, the interval widens.
For example, when $\hat{p}=0.0001$, the lower bound of the 95% Wilson interval is $\approx 1.77\times 10^{-5}$. This implies that about 56,64956{,}649 repetitions ($\approx 1/1.77\times 10^{5}$) may be required to detect such a flaky test with 95% confidence.
Not all flaky tests are this elusive. For instance, with $\hat{p}=0.01$, the lower bound of the 95% Wilson interval is $\approx 0.0121$, suggesting that about $82$ repetitions ($\approx 1/0.0121$) would be sufficient to detect failures with 95% confidence.
From the upper bound of the Wilson interval in Figure 2, we observe that for $\hat{p}=10^{-4}$, the smallest detectable value of $p$ is $4.5\times 10^{-4}$ at the 90% confidence level, $5.7\times 10^{-4}$ at 95%, and $8.5\times 10^{-4}$ at 99%.
Figure 3 illustrates how the required number of repetitions increases as $\hat{p}$ decreases. At moderate failure rates (e.g. $\hat{p}\approx 10^{-2}$), all confidence levels require a few hundred repetitions. As $\hat{p}$ approaches $10^{-4}$, the number of required repetitions rises sharply into the tens of thousands, and the gap between confidence levels widens. This highlights the high cost of detecting extremely rare flaky behaviors with strong statistical confidence.
Figure 3: Number of repetitions required to rule out extremely small true failure probabilities using Wilson lower confidence bounds. The $x$-axis shows the empirical estimate $\hat{p}$, and the $y$-axis shows the number of shots needed, computed as $1/L$, where $L$ is the Wilson lower bound at confidence levels 90%, 95%, and 99%. Higher confidence levels require more repetitions.

### v2

@@ 4 Study Results
Table 2: Test suite size and flakiness across terra releases. For each release, we report the total number of test cases, the number of non-skipped tests, the number of flaky tests (i.e., those that both pass and fail across 10,000 runs), the flakiness rate, and the contribution of that release to the global set of unique flaky tests. The flakiness rate is computed as the ratio of flaky test count over total test count. The share of all unique flaky tests is computed as the ratio of the flaky test count for a given release over the total number of unique flaky tests across all 290 releases. The final two rows aggregate totals across all releases.
@@ 4.3.2 Theoretical analysis
Figure 2: Wilson confidence intervals for the true failure probability $p$ as a function of the empirical estimate $\hat{p}$ when $N=10000$. For each confidence level $\gamma\in\{90\%,95\%,99\%\}$, the upper and lower bounds are shown as paired lines with matching color and line type. The reference line $p=\hat{p}$ is shown with a dot-dash pattern.
Figure 2 shows how the Wilson interval varies with $\hat{p}$ at different confidence levels $\gamma$. As $\hat{p}$ decreases, the interval widens.
We next connect these bounds to a conservative rerun budget for flakiness detection. Let $X$ be the number of observed failures after $n$ executions, so that $X\sim\mathrm{Binomial}(n,p)$. In general, the probability of observing exactly $k$ failures is
$\Pr(X=k)=\binom{n}{k}p^{k}(1-p)^{n-k},\qquad k=0,1,\dots,n.$
Therefore, the probability of detecting flakiness by observing at least one failure is
$\Pr(X\geq 1)=1-\Pr(X=0)=1-\binom{n}{0}p^{0}(1-p)^{n}=1-(1-p)^{n}.$ (1)
Given a target detection probability $q\in(0,1)$, we require
$1-(1-p)^{n}\geq q\quad\Leftrightarrow\quad(1-p)^{n}\leq 1-q\quad\Rightarrow\quad n\geq\frac{\ln(1-q)}{\ln(1-p)},$ (2)
as $\ln(1-p)<0$.
In practice, $p$ is unknown. Let $[L,U]$ be the Wilson confidence interval for $p$ with nominal coverage $\gamma$, computed from $N$ runs, and let $L=L_{\gamma}(\hat{p},N)$ denote its lower bound. Using $L$ as a conservative lower bound on the unknown $p$, and noting that Equation 1 is monotonically increasing in $p$, the worst case occurs at $p=L$. Therefore, a conservative rerun budget that targets detection probability at least $q$ is
$n_{q,\gamma}(\hat{p},N)=\left\lceil\frac{\ln(1-q)}{\ln\left[1-L_{\gamma}(\hat{p},N)\right]}\right\rceil,$
where $\lceil\cdot\rceil$ denotes the ceiling function.
Figure 3 plots $n_{q,\gamma}(\hat{p},N)$ versus $\hat{p}$ (for fixed $N$ and $q$), illustrating how required repetitions grow rapidly as $\hat{p}$ decreases and increase further with higher confidence $\gamma$. For example, with $q=0.95$ and $\gamma=0.95$, when $N=10000$ and $\hat{p}=10^{-4}$, then the 95% Wilson lower bound is $L_{0.95}\approx 1.77\times 10^{-5}$. Plugging these values into Equation 2, yields $n_{0.95,0.95}(0.0001,10000)\approx 1.69\times 10^{5}$ repetitions. When $\hat{p}=0.01$ and $L\approx 8.23\times 10^{-3}$, this reduces to $n_{0.95,0.95}(0.01,10000)\approx 363$ repetitions.
Figure 3: Conservative rerun budget for flakiness detection using Wilson lower confidence bounds. The $x$-axis shows the empirical estimate $\hat{p}$ from $N=10000$ observed runs, and the $y$-axis shows the required number of repetitions $n_{q,\gamma}(\hat{p},N)$ with $q=95\%$ and $\gamma\in\{90\%,95\%,99\%\}$. Higher confidence levels require more repetitions.